OpenStudy (anonymous):

what times what equals 1536 but when u add/subtract them it eguals -16

OpenStudy (anonymous):

Set up 2 equations: x*y=1536 x+y=-16 (based on your wording?) You know that x = 1536/y, plug it into the second equation and solve for y. Then take that value and get the value for x.

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

what are the two numbers

OpenStudy (anonymous):

x and y are just what we call the first and second numbers that you're working with. When you plug that new value of x into the second equation, you get 1536/y + y = -16. You can now multiply every term by y to eliminate the fraction, and solve the quadratic to get the values of y.

OpenStudy (anonymous):

i still dont get it? i never learned that way?^ can you please just tell me the two number and i can work the rest of the problem on from there

OpenStudy (anonymous):

u need to use x-y=-6 because if u use x+y=-16, the quadratic u get wont have real solution

OpenStudy (anonymous):

can someone please just tell me the two numbers